Solve the system of equations below using elimination by addition. What are the solutions?
4m-n=26
m-4n=-16 Answer by mananth(16949) 
multiply (1)by -4
Multiply (2) by 1
-16 m 4 n = -104
1 m -4 n = -16
Add the two equations
-15 m = -120
/ -15
m = 8
plug value of m in (1)
4 m + -1 n = 26
32 + -1 n = 26
-1 n = -6
n = 6
Ans m = 8
n = 6
(8,6) is the solution